Solve the problems below and show the complete solution.
Determine the pH of a solution consisting of 0.15 M NH4Cl and 1.5 M NH3. The Kb of NH3 is 1.8x10-5.
1
Expert's answer
2020-12-06T13:53:49-0500

"NH4Cl=0.15m"

"NH3=1.5m"

"Kb=1.8\u00d710-5"

"pKb=-logKb"

"pKb=-log(1.8\u00d710-5)"

"pKb=4.745"

"POH=pKb+log(salt\/acid)"

"POH=4.745+log(0.15\/1.5)"

"POH=3.745"

"PH+POH=14"

"PH=14-3.745"

"PH=10.26"